# Relative uniform convergence and Archimedean property in pre-ordered vector spaces

### v1

@@ Abstract
It is proved that, for a pre-ordered vector space $X$, the quotient space $(X/A,[W])$ is the Archimedeanization of $X$, where $W$ is the closure of the positive wedge $X_{+}$ in the ru-topology, $A=W\cap(-W)$, and $[W]$ is the quotient set of $W$ in $X/A$.
@@ 1 Introduction
The present paper is devoted to an interplay between the Archimedean property and relatively uniform convergence in (pre-)ordered vector spaces. In Theorem 3.3, we prove that the quotient set $[W]$ of the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a pre-ordered vector space $X$ is an Archimedean cone in quotient space $X/(W\cap(-W))$. The main result of our paper is Theorem 3.5 which establishes that $(X/(W\cap(-W)),[W])$ is the Archimedeanization of $X$.
@@ 2 Preliminaries
A subset $W$ of a vector space $X$ is a wedge if $W+W\subseteq W\ \ \text{\rm and}\ \ tW\subseteq W\ \ \text{\rm for}\ \ t\geqslant 0$. A wedge $C$ is a cone whenever $C\cap(-C)=\{0\}$. Under an ordered vector space (briefly, OVS) we understand a vector space $X$ together with a cone (denoted by $X_{+}$ and called a positive cone of $X$). Replacing a cone by a wedge, we also consider a pre-ordered vector space (briefly, POVS). A wedge $W$ in a vector space $X$ is majorizing (or generating) if
[…]
for every $x\in X$.
Every OVS (POVS) $X$ is equipped with a partial (pre-)order
@@
Every pair of vectors $\{a,b\}$ in an OVS (POVS) $X$ produces an order interval (possibly empty, if $a\not\leqslant b$):
@@
Clearly, each almost Archimedean POVS must be an OVS. Moreover, every Archimedean OVS is almost Archimedean. The converse is not true even for two-dimensional OVSs (see Example 2.5 in [3]). It is also straightforward to see: a POVS $X$ is almost Archimedean iff $X_{+}$ does not contain a straight line.
A net $(x_{\alpha})$ in a POVS $X$ ru-converges to $x\in X$ (shortly, $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x$) if, for some (a regulator of the convergence) $w\in X_{+}$, there exists a sequence $(\alpha_{n})$ of indices with $x_{\alpha}-x\leqslant\pm\frac{1}{n}w$ for $\alpha\geqslant\alpha_{n}$. Whenever we need to specify a regulator $w$, we write $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x(w)$.
Clearly, every ru-convergent net in an OVS $X$ has a unique ru-limit iff $X$ is almost Archimedean. As the ru-convergence is sequential, we can always restrict ourselves to ru-convergent sequences. The ru-convergence is an abstraction of the classical uniform convergence of functions in $C[0,1]$. The notion of ru-convergence was introduced by L.V. Kantorovich [9].
3 Closedness of the positive wedge in ru-topology vs Archimedean property in pre-ordered vector spaces
Let $X/A$ be a quotient space. A quotient set of $B\subseteq X$ is the set $B]=\{[x]\in X/A:x\in B\}$ of all equivalence classes formed from a set $B$ in the quotient space $X/A$.
The following lemma is a standard fact.
@@ 3 Closedness of the positive wedge in ru-topology vs Archimedean property in pre-ordered vector spaces
Let $W$ be a wedge in a vector space $X$. Then $[W]$ is a cone in $X/A$, where $A=W\cap(-W)$. Moreover, $[W]$ is majorizing in $X/A$ whenever $W$ is majorizing in $X$.
@@
Clearly, $A$ is a subspace of $X$ and $[W]$ is a wedge in $X/A$. To see that $[W]$ is a cone it suffices to show $W]\cap-[W]\subseteq A$. Let $x]\in[W]\cap-[W]$. Then $x=x_{1}+A=-x_{2}+A$ for some $x_{1},x_{2}\in W$. It follows $x\in(W+W)\cap(-W-W)\subseteq W\cap(-W)=A$, and hence $[x]=0$. The rest is trivial. ∎
An OVS $X$ is a vector lattice if $x^{+}:=\inf\bigl\{X_{+}\cap(x+X_{+})\bigl\}$ exists for every $x\in X$. It is easy to see that every almost Archimedean vector lattice is Archimedean. An order ideal $A$ in a vector lattice $X$ will be referred to as a lattice ideal whenever it satisfies the property: $a\in A\Longrightarrow|a|\in A$.
W.A.J. Luxemburg and L.C-Jr. Moore proved in [10] that the ru-closed subsets of a vector lattice $X$ are exactly the closed sets of the so-called ru-topology on $X$. This topology can be also introduced on an arbitrary POVS (cf., [8]).
@@
It follows immediately from Definition 3.1 that the ru-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ with the property $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
The ru-topology is not linear in general. It is proved in [10, Theorem 4.2] that every lattice homomorphism between vector lattices is continuous in the ru-topology. Indeed, it is true for arbitrary positive operators between POVSs, and if additionally the positive wedge in the co-domain is majorizing, it is also true for every order bounded operator.
@@ 1. $i)$
    If $Y_{+}$ is majorizing and $T$ is order bounded then $T$ is $\tau_{ru}$-continuous.
@@
Let $S\subseteq Y$ $\tau_{ru}$-closed. It suffices to show that $T^{-1}(S)$ is $\tau_{ru}$-closed. To this end, let $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$. By passing to a subsequence, we may suppose $x_{n}-x\leqslant\pm\frac{1}{n}u$ for some $u\in X_{+}$ and all $n\in\mathbb{N}$. By the definition of $\tau_{ru}$-topology, it remains to show $x\in T^{-1}(S)$.
If $T\geqslant 0$ then $Tx_{n}-Tx\leqslant\pm\frac{1}{n}Tu$, or $Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$. Therefore, $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$ and, since $S$ is $\tau_{ru}$-closed, we infer $Tx\in S$, and hence $x\in T^{-1}(S)$.
If $Y_{+}$ is majorizing and $T$ is order bounded, then $T$ is ru-continuous by [5, Theorem 2.4 ii)] As $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ then $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$. Thus, $Tx\in S$, and again $x\in T^{-1}(S)$. ∎
It was observed by T. Ito [6] that a vector lattice $X$ is Archimedean iff $X_{+}$ is ru-closed. We show that it is happened also with every POVS.
Lemma 3.3.
A POVS $X$ with a positive wedge $X_{+}$ is Archimedean iff $X_{+}$ is ru-closed.
@@
Lemma 3.3 should be compared with the following classical fact (cf., [1, Lemma 2.3]).
@@
If a POVS $X$ admits a linear topology $\tau$ for which its wedge $X_{+}$ is $\tau$-closed, then $X$ is Archimedean.
The $\tau_{ru}$-topology is not necessarily linear. However, as it was mentioned above, the $\tau_{ru}$-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$. Therefore, we obtain the following corollary of Lemma 3.3.
Corollary 3.1.
A POVS $X$ is Archimedean iff $X_{+}$ is $\tau$-closed in some topology $\tau$ on $X$ possessing the property: $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
@@
The forward implication of Theorem 3.2 is almost trivial in the OVS case [3, Proposition 3.2(a)]. We do not have yet an extension of Theorem 3.2 to any reasonable class of POVS beside vector lattices. However, for a very natural ru-closed order ideal $A$ in an arbitrary POVS $X$ the quotient $X/A$ turns to be Archimedean.
@@
Let $X$ be a POVS with a positive wedge $X_{+}$. Then $[W]$ is an Archimedean cone in $X/A$, where $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap(-W)$. Moreover, the cone $[W]$ is majorizing whenever $X_{+}$ is majorizing.
@@
Clearly, $A$ is an order ideal in $X$. By Lemma 3.1, $[W]$ is a cone in $X/A$. The set $W$ (and, also W-W) is $\tau_{ru}$-closed. Let $n[x]\leqslant[y]$ for all $n\in\mathbb{N}$ and some $x\in X,y\in W$. Then $x]\leqslant\frac{1}{n}[y]$, and hence $x-\frac{1}{n}y]\in-[W]$ for all $n\in\mathbb{N}$. So, there exists $(a_{n})_{n}$ in $A$ such that $x-\frac{1}{n}y+a_{n}\in-W$ for all $n\in\mathbb{N}$. Note that $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ (with a regulator 0). As W-W is $\tau_{ru}$-closed, we infer from $\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}0$ and $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ that $x-\frac{1}{n}y+a_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$, and hence $x\in-W$. It follows $x]\in-[W]$, or $x]\leqslant 0$. So, $(X/A,[W])$ is Archimedean. Due to Lemma 3.1, $[W]$ is majorizing whenever $X_{+}\subseteq W$ is majorizing. ∎
It was proved in [12] that every OVS with an order unit has a unique (up to an order isomorphism) Archimedization. This result was extend further to arbitrary OVS in [3]. The following definition was introduced in [3, Definition 5].
@@
By [3, Corollary 3.5], any almost Archimedean cone can be embedded into an Archimedean cone. In particular, the Archimedeanization of any almost Archimedean OVS space $X$ can be obtained via replacing $X_{+}$ by the intersection of all Archimedean cones containing $X_{+}$. The following result is proved in [3, Theorem 1].
Proposition 3.4.
Every OVS has the Archimedeanization.
We present a construction of the Archimedeanization (based on Theorem 3.3), which seems to be more elementary than one given in [3].
@@
Let $X$ be a POVS with a positive wedge $X_{+}$. Let $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap(-W)$. Then $(X/A,[W])$ is the Archimedeanization of $X$.

### v3

@@ Abstract
It is proved that, for a pre-ordered vector space $X$, the quotient space $(X/A,[W])$ is the Archimedeanization of $X$, where $W$ is the closure of the positive wedge $X_{+}$ in the ru-topology, $A=W\cap-W$, and $[W]$ is the quotient set of $W$ in $X/A$.
@@ 1 Introduction
The present paper is devoted to an interplay between the Archimedean property and relatively uniform convergence in (pre-)ordered vector spaces. In Theorem 3.3, we prove that the quotient set $[W]$ of the $\tau_{ru}$-closure of the positive wedge $X_{+}$ of a pre-ordered vector space $X$ is an Archimedean cone in quotient space $X/(W\cap-W)$. The main result of our paper is Theorem 3.4 which establishes that $(X/(W\cap-W),[W])$ is the Archimedeanization of $X$.
@@ 2 Preliminaries
A subset $W$ of a vector space $X$ is a wedge if $W+W\subseteq W\ \ \text{\rm and}\ \ tW\subseteq W\ \ \text{\rm for}\ \ t\geqslant 0$. A wedge $C$ is a cone whenever $C\cap-C=\{0\}$. Under an ordered vector space (briefly, OVS) we understand a vector space $X$ together with a cone (denoted by $X_{+}$ and called a positive cone of $X$). Replacing a cone by a wedge, we also consider a pre-ordered vector space (briefly, POVS). A wedge $W$ in a vector space $X$ is generating if WW=XW-W=X. Every OVS (POVS) $X$ is equipped with a partial (pre-)order
@@
Each pair of vectors $\{a,b\}$ in an OVS (POVS) $X$ produces an order interval (possibly empty, if $a\not\leqslant b$):
@@
Clearly, every almost Archimedean wedge is a cone and every Archimedean cone is almost Archimedean. On the other hand, even a two-dimensional almost Archimedean OVS need not to be Archimedean (see Example 2.5 in [3]). Furthermore, an OVS $X$ is almost Archimedean if and only if $X_{+}$ does not contain a straight line.
A net $(x_{\alpha})$ in a POVS $X$ ru-converges to $x\in X$ (shortly, $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x$) if, for some (regulator of the convergence) $w\in X_{+}$, there exists a sequence $(\alpha_{n})$ of indices with $\pm(x_{\alpha}-x)\leqslant\frac{1}{n}w$ for $\alpha\geqslant\alpha_{n}$. Whenever we need to specify a regulator $w$, we write $x_{\alpha}\operatorname{\xrightarrow[]{{\it ru}}}x(w)$.
Clearly, every ru-convergent net in an OVS $X$ has a unique ru-limit iff $X$ is almost Archimedean. As the ru-convergence is sequential, we can always restrict ourselves to ru-convergent sequences. The ru-convergence is an abstraction of the classical uniform convergence of functions in $C[0,1]$. The notion of ru-convergence was introduced by L.V. Kantorovich in [10].
3 Closedness of a positive wedge in ru-topology vs Archimedean property
Let $X/A$ be a quotient space. A quotient set of $B\subseteq X$ is the set $B]=\{[x]\in X/A:x\in B\}$ of all equivalence classes formed from a set $B$ in the quotient space $X/A$. The following lemma is a standard fact.
@@ 3 Closedness of a positive wedge in ru-topology vs Archimedean property
Let $W$ be a wedge in a vector space $X$. Then $[W]$ is a cone in $X/A$, where $A=W\cap-W$. Moreover, $[W]$ is generating in $X/A$ if and only if $W$ is generating in $X$.
@@
Clearly, $A$ is a subspace of $X$ and $[W]$ is a wedge in $X/A$. To see that $[W]$ is a cone it suffices to show $W]\cap-[W]\subseteq A$. Let $x]\in[W]\cap-[W]$. Then $x=x_{1}+A=-x_{2}+A$ for some $x_{1},x_{2}\in W$. It follows $x\in(W+W)\cap(-W-W)\subseteq W\cap-W=A$, and hence $[x]=0$.
Suppose WW=XW-W=X and let $x]\in X/A$. Then, $x=y_{1}-y_{2}$ for some $y_{1},y_{2}\in W$, and hence $x]=[y_{1}-y_{2}]=[y_{1}]-[y_{2}]\in[W]-[W]$. Thus, $[W]$ is generating in $X/A$.
Suppose [W][W]=X/A[W]-[W]=X/A and let $x\in X$. Take $y_{1},y_{2}\in W$ with $x]=[y_{1}]-[y_{2}]=[y_{1}-y_{2}]$. So, for some $z\in A$ we have $x+z=y_{1}-y_{2}$, and hence $x=y_{1}-(y_{2}+z)\in W-W$. Therefore, $W$ is generating in $X$. ∎
An OVS $X$ is a vector lattice if $x^{+}:=\sup\{x,0\}$ exists for every $x\in X$. It is easy to see that every almost Archimedean vector lattice is Archimedean. An order ideal $A$ in a vector lattice $X$ will be referred to as a lattice ideal whenever it satisfies the property: $a\in A\Longrightarrow|a|\in A$.
W.A.J. Luxemburg and L.C-Jr. Moore proved in [11] that the ru-closed subsets of a vector lattice $X$ are exactly the closed sets of the so-called ru-topology on $X$. Such a topology can be also introduced on an arbitrary POVS [9].
@@
It follows immediately from Definition 3.1 that the ru-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ with the property $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$. Clearly, $\tau_{ru}$ is translation invariant. In general, the topology $\tau_{ru}$ is not linear.
It is useful to describe constructively the $\tau_{ru}$-closure of a subset of a POVS. The idea is coming from [11], where the pseudo uniform closure of a subset of a vector lattice was studied. Adopting the idea to the POVS setting, for a subset $S$ of a POVS $X$ we denote $S^{(0)}_{ru}:=S$; $S^{(1)}_{ru}:=S^{\prime}_{ru}$ the set of all $x\in X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ for some sequence $(x_{n})_{n}$ in $S$; $S^{(\alpha)}_{ru}:=\bigl(S^{(\alpha-1)}_{ru}\bigl)^{\prime}_{ru}$ when the ordinal $\alpha$ is not limit; and $S^{(\alpha)}_{ru}:=\bigcup\limits_{\beta\in\alpha}S^{(\beta)}_{ru}$ when $\alpha$ is limit.
Lemma 3.2.
Let $S$ be a subset of a POVS $X$. Then:
  1. $i)$
    $S^{(\alpha_{1})}_{ru}\subseteq S^{(\alpha_{2})}_{ru}$ whenever $\alpha_{1}\leqslant\alpha_{2}$.
  2. ii)ii)
    $\overline{S}_{ru}=\bigcup\limits_{\alpha\in\omega_{1}}S^{(\alpha)}_{ru}$, where $\omega_{1}$ is the first uncountable ordinal.
  3. $iii)$
    $S\subseteq X_{+}-X_{+}\Longleftrightarrow\overline{S}_{ru}\subseteq X_{+}-X_{+}$.
Proof.
$i)$ It is straightforward.
ii)ii) It is similar to the proof of [11, Theorem 3.3], where the vector lattice case is considered.
$iii)$ It suffices to prove $S\subseteq X_{+}-X_{+}\Longrightarrow S^{\prime}_{ru}\subseteq X_{+}-X_{+}$. So, let $S\subseteq X_{+}-X_{+}$ and $S\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$. By passing to a subsequence, if necessary, we may assume that $\pm(x_{n}-x)\leqslant\frac{1}{n}w$ for some $w\in X_{+}$ and all $n\in\mathbb{N}$. In particular, $x-x_{1}\leqslant w$, or $x\leqslant x_{1}+w$. Since $x_{1}=x_{1}^{1}-x_{1}^{2}$ for some $x_{1}^{1},x_{1}^{2}\in X_{+}$, we obtain $x\leqslant w+x_{1}^{1}$ and $x=(w+x_{1}^{1})-(w+x_{1}^{1}-x)\in X_{+}-X_{+}$. ∎
Lemma 3.3.
Let $D$ be a wedge in a POVS $X$. Then:
  1. $i)$
    $D^{(\alpha)}_{ru}$ is a wedge in $X$ for every ordinal $\alpha$.
  2. ii)ii)
    $\overline{D}_{ru}$ is a wedge in $X$.
Proof.
$i)$ Let $\alpha$ be an ordinal. We have to prove that $t\cdot D^{(\alpha)}_{ru}\subseteq D^{(\alpha)}_{ru}$ for all $t\in\mathbb{R}_{+}$, and that $D^{(\alpha)}_{ru}+D^{(\alpha)}_{ru}\subseteq D^{(\alpha)}_{ru}$.
By definition, $D^{(0)}_{ru}=D$, so $D^{(0)}_{ru}$ is a wedge. Suppose $D^{(\beta)}_{ru}$ is a wedge for every $\beta<\alpha$. First, consider the case when $\alpha=\beta+1$ for some $\beta$. So, $D^{(\beta)}_{ru}$ is a wedge and $D^{(\alpha)}_{ru}=\bigl(D^{(\beta)}_{ru}\bigl)^{\prime}_{ru}$. Let $t\geqslant 0$ and $x\in D^{(\alpha)}_{ru}$. Find a sequence $(x_{n})$ in $D^{(\beta)}_{ru}$ with $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$. By passing to a subsequence, we may assume that $\pm(x_{n}-x)\leqslant\frac{1}{n}w$ for some $w\in X_{+}$ and all $n\in\mathbb{N}$. Then, $\pm(tx_{n}-tx)\leqslant\frac{1}{n}tw$ for all $n\in\mathbb{N}$. We conclude $D^{(\beta)}_{ru}\ni tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}tx$, and hence $tx\in\bigl(D^{(\beta)}_{ru}\bigl)^{\prime}_{ru}=D^{(\alpha)}_{ru}$. Now, let $\alpha$ be a limit ordinal, $t\geqslant 0$, and $x\in D^{(\alpha)}_{ru}=\bigcup\limits_{\beta\in\alpha}D^{(\beta)}_{ru}$. Then, $x\in D^{(\beta)}_{ru}$ for some $\beta<\alpha$. By the assumption, $tx\in tD^{(\beta)}_{ru}\subseteq D^{(\beta)}_{ru}\subseteq D^{(\alpha)}_{ru}$. We conclude again $tx\in D^{(\alpha)}_{ru}$.
The proof of $D^{(\alpha)}_{ru}+D^{(\alpha)}_{ru}\subseteq D^{(\alpha)}_{ru}$ is similar, we point out only the case of $\alpha=\beta+1$. So, let $x,y\in D^{(\beta+1)}_{ru}$, say $D^{(\beta)}_{ru}\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ and $D^{(\beta)}_{ru}\ni y_{n}\operatorname{\xrightarrow[]{{\it ru}}}y$. By passing to subsequences, we may assume $\pm(x_{n}-x)\leqslant\frac{1}{n}w$ and $\pm(y_{n}-y)\leqslant\frac{1}{n}v$ for some $w,v\in X_{+}$ and all $n\in\mathbb{N}$. Then, $\pm((x_{n}+y_{n})-(x+y))\leqslant\frac{1}{n}(w+v)$ for all $n\in\mathbb{N}$. As $D^{(\beta)}_{ru}$ is a wedge, we conclude $D^{(\beta)}_{ru}\ni(x_{n}+y_{n})\operatorname{\xrightarrow[]{{\it ru}}}(x+y)$, and hence $x+y\in\bigl(D^{(\beta)}_{ru}\bigl)^{\prime}_{ru}=D^{(\beta+1)}_{ru}$.
ii)ii) It follows from $i)$ due to Lemma 3.2. ∎
It is proved in [11, Theorem 4.2] that every lattice homomorphism between vector lattices is continuous in the ru-topology. Indeed, it is true for arbitrary positive operators between POVSs, and if additionally the positive wedge in the co-domain is generating, it is also true for every order bounded operator.
@@ 1. $i)$
    If $Y_{+}$ is generating and $T$ is order bounded then $T$ is $\tau_{ru}$-continuous.
@@
Let $S\subseteq Y$ $\tau_{ru}$-closed. It suffices to show that $T^{-1}(S)$ is $\tau_{ru}$-closed. To this end, let $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$. By passing to a subsequence, we may suppose $\pm(x_{n}-x)\leqslant\frac{1}{n}u$ for some $u\in X_{+}$ and all $n\in\mathbb{N}$. By the definition of $\tau_{ru}$-topology, it remains to show $x\in T^{-1}(S)$.
If $T\geqslant 0$ then $\pm(Tx_{n}-Tx)\leqslant\frac{1}{n}Tu$, or $Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$. Therefore, $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$ and, since $S$ is $\tau_{ru}$-closed, we infer $Tx\in S$, and hence $x\in T^{-1}(S)$.
If $Y_{+}$ is generating and $T$ is order bounded, then $T$ is ru-continuous by [5, Theorem 2.4 ii)]. Therefore, $T^{-1}(S)\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x$ implies $S\ni Tx_{n}\operatorname{\xrightarrow[]{{\it ru}}}Tx$. Thus, $S\ni Tx_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}Tx$. So, $Tx\in S$, and again $x\in T^{-1}(S)$. ∎
It was observed by T. Ito [7] that a vector lattice $X$ is Archimedean iff $X_{+}$ is ru-closed. We show that it is happened also in every POVS.
Proposition 3.1.
A POVS $X$ is Archimedean if and only if its positive wedge $X_{+}$ is ru-closed.
@@
Proposition 3.1 should be compared with the following classical fact (cf., [1, Lemma 2.3]).
Assertion 3.1.
Every POVS $X$ that admits a linear topology $\tau$ for which $X_{+}$ is $\tau$-closed is Archimedean.
It is worth mentioning a recent paper [6] by A.E. Gutman and I.A. Emelianenkov, where an exhaustive description of the class of locally convex spaces in which all Archimedean cones are topologically closed is obtained. However, the $\tau_{ru}$-topology is not necessarily linear. But, as it was mentioned above, $\tau_{ru}$-topology on a POVS $X$ is the strongest topology $\tau$ on $X$ such that $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$. This leads to the following consequence of Proposition 3.1.
@@
Let $X$ be a POVS. The following assertions are equivalent.
  1. $i)$
    $X$ is Archimedean.
  2. ii)ii)
    $X_{+}$ is ru-closed.
  3. $iii)$
    $X_{+}$ is closed in the topology $\tau_{ru}$.
  4. $iv)$
    $X_{+}$ is closed in some topology $\tau$ on $X$ possessing the property: $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$.
Proof.
Equivalence $i)\Longleftrightarrow ii)$ is the statement of Proposition 3.1, $ii)\Longrightarrow iii)$ follows from Definition 3.1, and $iii)\Longrightarrow iv)$ is trivial.
$iv)\Longrightarrow ii)$ Let $X_{+}$ be closed in a topology $\tau$ on $X$ satisfying $x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\Longrightarrow x_{n}\operatorname{\xrightarrow[]{\tau}}x$. If $X_{+}$ is not ru-closed, for some sequence $(x_{n})$ we have $X_{+}\ni x_{n}\operatorname{\xrightarrow[]{{\it ru}}}x\notin X_{+}$, and hence $X_{+}\ni x_{n}\operatorname{\xrightarrow[]{\tau}}x$. Since $X_{+}$ is $\tau$-closed, we obtain $x\in X_{+}$, a contradiction. ∎
@@ 4. $iv)$
The forward implication of Assertion 3.2 is almost trivial in the OVS case [3, Proposition 3.2(a)]. On the other hand, for an OVS $V$ from [3, Example 2.5], $V=V/\{0\}$ is not Archimedean although the ideal $\{0\}$ is ru-closed. We do not have yet an extension of Assertion 3.2 to reasonable classes of POVSs beside vector lattices. However, for a very natural ru-closed order ideal $A$ in an arbitrary POVS $X$ the quotient $X/A$ turns to be Archimedean.
@@
Let $X$ be a POVS with a positive wedge $X_{+}$. Then $[W]$ is an Archimedean cone in $X/A$, where $W$ is the $\tau_{ru}$-closure of $X_{+}$ and $A=W\cap-W$. Moreover, the cone $[W]$ is generating whenever $X_{+}$ is generating.
@@
By Lemma 3.3, $W$ is a wedge. Lemma 3.1 implies that $[W]$ is a cone in $X/A$. The set $W$ (and, also W-W) is $\tau_{ru}$-closed. Let $n[x]\leqslant[y]$ for all $n\in\mathbb{N}$ and some $x\in X,y\in W$. Then $x]\leqslant\frac{1}{n}[y]$, and hence $x-\frac{1}{n}y]\in-[W]$ for all $n\in\mathbb{N}$. So, there exists $(a_{n})_{n}$ in $A$ such that $x-\frac{1}{n}y+a_{n}\in-W$ for all $n\in\mathbb{N}$. Clearly, $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ (with regulator 0). By Lemma 3.2, $y\in X_{+}-X_{+}$. Then, $\pm y\leqslant u$ for some $u\in X_{+}$, and hence $\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}0$ (with regulator $u$). So, $x-\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}x$. As W-W is $\tau_{ru}$-closed, we infer from $x-\frac{1}{n}y\operatorname{\xrightarrow[]{{\it ru}}}x$ and $a_{n}\operatorname{\xrightarrow[]{{\it ru}}}0$ that $x-\frac{1}{n}y+a_{n}\operatorname{\xrightarrow[]{\tau_{{\it ru}}}}x$, and hence $x\in-W$. It follows $x]\in-[W]$, or $x]\leqslant 0$. Therefore, $(X/A,[W])$ is Archimedean.
Suppose $X_{+}$ is generating. Since $X_{+}\subseteq W$ then $W$ is generating too, and hence $[W]$ is generating by Lemma 3.1. ∎
It was proved in [13] that every OVS with an order unit has a unique (up to an order isomorphism) Archimedization. This result has being extend to arbitrary OVS in [3]. The next notion was introduced in [3, Definition 5].
@@
The following result is proved in [3].
Assertion 3.3.
([3, Theorem 1]) Every OVS has the Archimedeanization.
We present a construction of the Archimedeanization (based on Theorem 3.3), which looks more elementary than the one given in [3].
@@
Let $X$ be a POVS with a positive wedge $X_{+}$, $W$ the $\tau_{ru}$-closure of $X_{+}$, and $A=W\cap-W$. Then $(X/A,[W])$ is the Archimedeanization of $X$.